# Reversible Deep Equilibrium Models

### v1

@@ 6 Conclusion
We have introduced a new class of reversible implicit model, RevDEQs. The model uses a reversible fixed point algorithm to compute the forward pass and performs exact gradient backpropagation in constant memory with respect to model depth. This construction demonstrates state-of-the-art performance while significantly reducing the large number of function evaluations previously required by implicit architectures.
@@ B.1 Language Modelling on Wikitext-103
The model architecture is given by equation (7). We use an embedding size of $d=768$, with $\mathbf{z}\in\mathbb{R}^{d}$ and $W_{QKV}\in\mathbb{R}^{3d\times d}$. The MLP has depth 2 with hidden size $4d$; the GELU [49] activation function is used. The number of attention heads was 12. The token sequence length was 448 with a batch size of 32.
@@
We use the Adam optimiser with a weight decay of $10^{-4}$ and an initial learning rate of $3\times 10^{-4}$ [50]. The learning rate schedule is a decay on plateau schedule; we half the learning rate if the training loss does not increase for 10 steps (accumulated over the past 200 steps). Training is performed for a maximum of 10 epochs.
In Table 1, the Transformer-XL and DEQ results are from [10]. In Table 2, the DEQ results are from [10] and the DEQ with Jacobian regularisation results are from [20].
@@ B.2 Image Classification on CIFAR-10
We use the Adam optimiser with a weight decay of $10^{-4}$ and an initial learning rate of $10^{-3}$ [50]. The learning rate schedule is a decay on plateau schedule; we half the learning rate if the training loss does not increase for 10 steps (accumulated over the past 200 steps). Training is performed for a maximum of $100$ epochs or if the learning rate drops below $10^{-6}$. The batch size was 64.
@@
After each RevDEQ block we apply downsampling; for scales $i=1,2,3$ we use the pre-activation downsampling block from [45] with BatchNorm. At the final scale, $i=4$, we downsample using a global average pooling operation. Finally, the linear classification layer is applied to compute the logits.
@@
In Table 3, the monDEQ results are from [21]; the pcDEQ results are from [22]; the DEQ, MDEQ and explicit ResNet results are from [12]. The test accuracy is shown as mean $\pm$ std over 3 repeats.

### v2

@@ 6 Conclusion
We have introduced a new class of reversible implicit model, RevDEQs. The model uses a reversible fixed point algorithm to compute the forward pass and performs exact gradient backpropagation in constant memory with respect to model depth. This construction demonstrates improved performance while significantly reducing the large number of function evaluations previously required by implicit architectures.
@@ B.1 Language Modelling on Wikitext-103
The model architecture is given by equation (7). We use an embedding size of $d=1024$, with $\mathbf{z}\in\mathbb{R}^{d}$ and $W_{QKV}\in\mathbb{R}^{3d\times d}$. The MLP has depth 2 with hidden size $4d$; the GELU [hendrycks2016gaussian] activation function is used. The number of attention heads was 8. The token sequence length was 512 with a batch size of 32.
@@
We use the Adam optimiser with a weight decay of $0.1$ and a peak learning rate of $10^{-3}$ [loshchilov2019decoupled]. A cosine learning rate decay schedule was used with linear warm-up for $4000$ steps. Dropout was applied to the RevDEQ block with probability $0.1$. Training was performed for a maximum of 40 epochs.
In Table 1, the Transformer-XL and DEQ results are from [bai2021stabilizing]. In Table 2, the DEQ results with and without Jacobian regularisation are from [bai2021stabilizing].
@@ B.2 Image Classification on CIFAR-10
We use the Adam optimiser with a weight decay of $10^{-4}$ and an initial learning rate of $10^{-3}$ [loshchilov2019decoupled]. The learning rate schedule is a decay on plateau schedule; we half the learning rate if the training loss does not increase for 10 steps (accumulated over the past 200 steps). Training is performed for a maximum of $100$ epochs or if the learning rate drops below $10^{-6}$. The batch size was 64.
@@
After each RevDEQ block we apply downsampling; for scales $i=1,2,3$ we use the pre-activation downsampling block from [he2016identity] with BatchNorm. At the final scale, $i=4$, we downsample using a global average pooling operation. Finally, the linear classification layer is applied to compute the logits.
@@
In Table 3, the monDEQ results are from [winston2020monotone]; the pcDEQ results are from [gabor2024positive]; the DEQ, MDEQ and explicit ResNet results are from [bai2020multiscale]. The test accuracy is shown as mean $\pm$ std over 3 repeats.